Find y dV, where E is the solid bounded by the parabolic cylinder z = xand the planes y = 0 and 2 = 15 – 3y E
Answer #1

I M SSS ฯdv tz E 2-15 y=o 2=X2 -2 15-3y y 5 + 2 al y=3 515 15-2 15 IE 2 y dy dz dx 2015 -515 2 o x=515 515 15 X=-15 15-2 .2=XJis I- - 22 [ x 6 - 153 3 (15) x + 3 (15)-x2] x TF || Jis گرد) - 2구 7. -ts - 15 + 3(15)] [ * s) -> (15+ (is 리는 C -) 27 커2 512